# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v4

@@ 13. Outlook
Acknowledgements The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [13] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper, and to Makoto Yamashita for interesting discussions on [42] and [100] during his visit at the math department of Rome, Sapienza, and interesting comments on our work. Finally she thanks Pavel Etingof for informing her of [31].

### v5

@@ Appendix A Historical remarks on rigidity in affine Lie algebra and vertex operator algebra categories
The author is grateful to Yi-Zhi Huang for several comments on the proof of modularity for fusion categories of vertex operator algebras, and the history of Finkelberg-Kazhdan-Lusztig theorem. She is also grateful to Marco V. Giannone for pointing out to her incompleteness of the proof of Lemma 7.10 in version 7 of [10] and to Hans Wenzl for kind email correspondence on the centraliser algebras for the spinor representations, and for references on duality for tensor powers of the fundamental representation of Drinfeld-Jimbo quantum groups for the Lie types $B$, $D$, $G_{2}$, leading to the present paper. She also thanks Makoto Yamashita for interesting discussions on [39] and [100] during his visit at the math department of Rome, Sapienza, and interesting comments on our work. Finally she thanks Pavel Etingof for informing her of [28].
The author thanks Sapienza University of Rome for financial support Finanziamento di Ateneo per la Ricerca Scientifica, years 2018; 2019; 2020; 2021, 2023.
Author’s Note. Our work was the subject of a severe cyber-attack in 2021, recognized by the judicial authority in Italy. The legal actions currently undergone at the Italian Procura della Repubblica (2021–2023) concern the unauthorized use of materials from that project in other publications. The papers Carpi et al, Commun. Math. Phys., 390 (2022), 169; Carpi et al, Commun. Math. Phys. 402 (2023), 169; Adamo et al: Commun. Math. Phys. 404, (2023), 1231, are legally involved in the related penal procedure.